# Ridge Regression for Paired Comparisons: A Tractable New Approach, with Application to Premier League Football

### v1

@@ 4.2 True strengths distributed as Student t
The results just described were obtained under the assumption of normally distributed strengths, as was used to construct the pairwise empirical Bayes method. Less benign settings for pairwise empirical Bayes are considered next, by drawing the strengths from Student’s $t$ distribution with $\nu$ degrees of freedom. The simulated strengths were scaled so that their precision continues to be $\lambda$,
$\mu_{i}\sim\sqrt{\frac{\nu-2}{\lambda\nu}}t_{\nu},\quad\nu>2,\;i=1,\ldots,p,$
with $t_{\nu}$ denoting a Student’s $t$-distributed variate with $\nu$ degrees of freedom. We have considered degrees of freedom $\nu>2$ because otherwise the variance would be undefined. Drawing the strengths from a $t$ distribution is of interest because it allows a small number of teams to be distinctly stronger or weaker than the rest. Figures S1 and S2 of the Supplementary Material display the average logarithmic scores for the same 120 scenarios considered previously, but with strengths simulated from Student’s $t$ distributions with $\nu=8$ and $\nu=3$ degrees of freedom. There is no perceptible difference between the $t_{8}$ and normal distributions, as far as predictive performance is concerned; and even in the more extreme scenario with $t_{3}$-distributed strengths, the differences are rather small. While such simulation evidence is inevitably limited, it does provide some reassurance that the pairwise empirical Bayes method is robust to excess kurtosis in the strengths. Departures from normality due instead to skewness need not be studied, because the Thurstone-Mosteller model is identified through pairwise differences of the strengths which have a symmetrical distribution by construction.
@@ 5 Premier League
Since season 1995–1996, the English Premier League has been a double round robin between twenty teams. The current format of the tournament therefore consists of 38 match-weeks. We consider here all match results from the 28 seasons between 1995 and 2023. For each season, we trained the paired comparison models with the first 10, 15, 20, 25 and 30 match-weeks and predicted the results of the rest of the tournament. The paired comparison models were estimated by maximum likelihood estimation using the polr function from the MASS package (Venables and Ripley, 2002), by bias-reduced maximum likelihood for the cumulative logistic regression model developed in Kosmidis (2014) and implemented in the R function bpolr available through the supplementary materials of that paper, and by our pairwise empirical Bayes method.
As a reference forecast, we considered the naive predictions calculated from long-term frequencies of home wins, draws and away wins pooled across all 28 seasons of the Premier League. Those seasons saw 46% home wins, 25% draws and 29% away wins out of a total number of $10,640$ matches. The logarithmic score for the naive forecast is thus
$\text{LS}_{\text{naive}}=-0.46\log(0.46)-0.25\log(0.25)-0.29\log(0.29)=1.06.$
@@
We now illustrate the above findings in more detail for the most recent Premier League season included in our study. Figure 3 displays the logarithmic skill scores for the 2022–23 season. When only the first ten Premier League matchweeks are used for training, the predictions calculated from maximum likelihood estimates are much worse even than the naive forecast, whereas the pairwise empirical Bayes method substantially out-performs the naive forecaster. The predictions calculated with bias reduced maximum likelihood clearly improve upon maximum likelihood, but still not enough to produce predictions with acceptable precision when calculated from only ten matchweeks. The performance of maximum likelihood estimation does improve as the number of matchweeks used for training increases, but in any case pairwise empirical Bayes remains superior to maximum likelihood with or without bias correction.
Figure 3: Logarithmic skill scores for Premier League 2022–2023. The predictions are computed with strengths estimated using maximum likelihood (MLE), bias-reduced maximum likelihood (BRMLE) and pairwise empirical Bayes (PEB). The reference method is a naive forecast that uses only long-term frequencies of home wins, draws and away wins and corresponds to a zero logarithmic skill score (grey solid line).
Further insight into the relative predictive performance of the three methods can be obtained from Figure 4, which shows estimated strength parameters for the twenty Premier League teams after ten weeks of matches and at the end of the season. The plots show clearly the strong shrinkage effect of the ridge estimator tuned by the pairwise empirical Bayes method, relative to maximum likelihood estimation which over-emphasises match results seen in the early stages of the tournament.
After ten weeks of the 2022–23 season, Arsenal were at the top of the table with 27 points (nine wins and one defeat), followed by Manchester City with 26 points (eight wins and two draws). With these results, the maximum likelihood estimated strengths were $3.22$ (Arsenal) and $2.50$ (Manchester City). At the end of the season the league champions were Manchester City with a five-point lead over Arsenal, with maximum likelihood estimated strengths $1.35$ (Arsenal) and $1.66$ (Manchester City). While bias-reduced maximum likelihood estimates after ten matchweeks are sensibly attenuated to $2.55$ (Arsenal) and $2.02$ (Manchester City), still they are substantially larger then the corresponding strength estimates at the end of the season.
The ridge-estimated strengths are instead quite stable from the beginning to the end of the season: the strength estimates obtained by pairwise empirical Bayes moved only from $0.64$ to $0.62$ for Arsenal and from $0.60$ to $0.75$ for Manchester City, for example, between the 10th and 38th matchweek. Thus over-confident prediction of future match outcomes, based on only a few match results early in the season, is avoided. This example illustrates well the general importance of shrinkage estimation, for prediction with good precision.
At the other end of the Premier League table the picture is similar. After ten matchweeks of the 2022–23 season, Leicester City had only 4 league points (from one win, one draw and eight defeats). Leicester’s corresponding strength estimate via maximum likelihood at that time was very low, at $-2.99$. By the end of the season, the maximum likelihood estimate of Leicester’s strength had improved substantially to $-0.86$. The ridge estimates of strength are again much more stable, with Leicester moving only from $-0.59$ after ten weeks to $-0.40$ at the end of the season. Leicester were relegated, still, but not by the large margin that would have been predicted after ten weeks based on maximum likelihood (with or without bias reduction). Ridge estimation via pairwise empirical Bayes, on the other hand, gave a much more realistic picture of Leicester City’s early-season standing.
Figure 4: Estimated strength parameters of the twenty teams in Premier League season 2022–2023 after ten matchweeks and at the end of the season (38 matchweeks). Strengths are estimated by the maximum likelihood (MLE), bias-reduced maximum likelihood (BRMLE) and pairwise empirical Bayes (PEB) methods, in all cases with their sum constrained to zero. Highlighted are the estimated strengths of Arsenal (ARS), Manchester City (MNC) and Leicester City (LEI).
@@ 6 Final Remarks
The main purposes of this paper were to discuss ridge estimation of paired-comparison models motivated by applications in sport tournaments, and to develop the simple and effective pairwise empirical Bayes method for tuning the ridge penalty in the Thurstone-Mosteller model. As expected, the shrinkage provided by a ridge penalty yields predictions with substantially higher precision than standard maximum likelihood estimation. Indeed when a paired-comparison model is trained with relatively few matches, as is the case for very sparse tournaments or in the early stages of a round robin or similar tournament, prediction based on maximum likelihood is actually worse than the naive predictor that uses only the the long-term proportions of wins, draws and losses.
The focus here was on ‘static’ models, with team strength assumed constant through each season. A potentially fruitful direction for future work is to explore penalized estimation for dynamic paired-comparison models, in which the strength parameters vary during the season. In earlier work Fahrmeir and Tutz (1994) and Knorr-Held (2000) discuss dynamic models where the strength parameters evolve according to random walks, whereas Cattelan, Varin and Firth (2013) describe serial dependence in the strengths through an exponential smoothing scheme. Other dynamic models for football have used the number of goals scored by the teams, or their difference, as the response variable, rather than the match outcome as was used in this work; an example of dynamic modelling of the number of goals scored is Koopman and Lit (2015).
The simulations of Section 4 and the application to English Premier League show that the amount of shrinkage needed for good predictive performance when the training set and the number of teams are both small is appreciably greater than the shrinkage inherent in the bias-reduced maximum likelihood method of Firth (1993). This is unsurprising: it is well known that unbiasedness typically is not the main consideration for predictive performance. Kosmidis and Firth (2021) suggest that using a suitably scaled-up version of the bias-reducing penalty might be used instead, to improve predictive performance at the cost of some additional bias. In sport tournaments such as those considered in this paper, an appropriate scale factor to use could perhaps be computed through ‘leave-one-matchweek-out’ cross-validation.
@@
Supplementary material provides the simulation results described in Section 4.2 where strengths are drawn from a Student $t$ distribution. The Premier League data and R code to reproduce the analysis and graphs of Section 5 are available in the public repository https://github.com/crisvarin/peb.

### v2

@@ 4.2 True strengths distributed as Student t
The results just described were obtained under the assumption of normally distributed strengths, the assumption used to construct the pairwise empirical Bayes method. Less benign settings for pairwise empirical Bayes are considered next, by drawing the strengths from Student’s $t$ distribution with $\nu$ degrees of freedom. The simulated strengths are scaled so that their precision continues to be $\lambda$, $\mu_{i}\sim t_{\nu}\sqrt{(\nu-2)/(\lambda\nu)}$, $\nu>2$, $i=1,\ldots,p,$ with $t_{\nu}$ denoting a Student’s $t$-distributed variate with $\nu$ degrees of freedom. We take $\nu>2$ since otherwise the variance would be undefined. Drawing the strengths from a $t$ distribution is of interest because it allows a small number of items to be distinctly stronger or weaker than the rest. Figures S1 and S2 of the Supplementary Material display the average logarithmic scores for the same 120 scenarios considered previously, but with strengths simulated from Student’s $t$ distributions with $\nu=8$ and $\nu=3$ degrees of freedom. As far as predictive performance is concerned, there is no perceptible difference between the $t_{8}$ and normal distributions; and even in the more extreme scenario with $t_{3}$-distributed strengths, the differences are rather small. While such simulation evidence is inevitably limited, it does provide some reassurance that the pairwise empirical Bayes method is robust to excess kurtosis in the strengths.
@@ 5 Premier League
As an illustration of the ridge method we consider estimation of the strength of football teams in a round-robin tournament, before the tournament is complete. Since season 1995–1996, the English Premier League has been a double round robin between twenty teams. The current format of the tournament therefore consists of 38 match-weeks. We consider here all the results of the $10,640$ matches of the 28 seasons between 1995 and 2023, in order to evaluate the robustness of our method over a long period characterized by substantial changes in (English) football. For each season, we trained the paired comparison models with the first 10, 15, 20, 25 and 30 match-weeks and predicted the results of the rest of the tournament.
The problem we consider in this illustration is different from sequential prediction of next week’s match results given the team’s past results, for which it would be more appropriate to consider dynamic paired comparison models, such as the state-space approaches of Knorr-Held (2000); Fahrmeir and Tutz (1994), the exponential smoothing of Cattelan et al. (2013), the time-weighted likelihood of McHale and Morton (2011), the kernel smoothing of Bong et al. (2020), or the recent nonparametric spectral ranker of Tian et al. (2024).
In our illustration, the Thurstone-Mosteller model with ties (which here will be called ‘draws’, the word used in football) was estimated by maximum likelihood estimation using the polr function from the MASS package (Venables and Ripley, 2002), by bias-reduced maximum likelihood for the cumulative logistic regression model developed in Kosmidis (2014) and implemented in the R function bpolr available through the supplementary materials of that paper, and by our pairwise empirical Bayes method using R code available at https://github.com/crisvarin/peb.
As a reference forecast, we considered the naive predictions calculated from long-term frequencies of home wins, draws and away wins pooled across all 28 seasons of the Premier League. Those seasons saw 46% home wins, 25% draws and 29% away wins out of a total number of $10,640$ matches. The logarithmic score for the naive forecast is thus $\text{LS}_{\text{naive}}=-0.46\log(0.46)-0.25\log(0.25)-0.29\log(0.29)=1.06.$
@@ 6 Final Remarks
The main purposes of this paper were to discuss ridge-penalized estimation of paired-comparison models, and to develop the simple and effective pairwise empirical Bayes method for tuning the ridge penalty in the Thurstone-Mosteller model. As expected, the shrinkage provided by a ridge penalty yields predictions with substantially higher precision than standard maximum likelihood estimation. Indeed when a paired-comparison model is trained with relatively few observations, prediction based on maximum likelihood is actually worse than the naive predictor that uses only the marginal proportions of (home win, away win, draw) outcomes. The poor predictive performance of maximum likelihood when the training set is small is due to overfitting; the shrinkage achieved through a ridge penalty, tuned by the pairwise empirical Bayes method, effectively eliminates that problem.
The simulations of Section 4 show that the amount of shrinkage needed for good predictive performance, especially when the training set and the number of items are both small, is appreciably greater than the shrinkage inherent in the bias-reduced maximum likelihood method of Firth (1993). This is unsurprising: it is well known that unbiasedness typically is not the main consideration for predictive performance.
The authors are not aware of other work focused on estimation of paired comparison models with a ridge penalty. Some previous work in team sports modeling has considered more traditional linear ridge penalty models to assess the contribution of individual players to a team’s performance. A recent example is the adjusted plus/minus regularized model discussed in Matano et al. (2023); see also the references therein.
@@
Supplementary material provides the simulation results described in Section 4.2 where strengths are drawn from a Student $t$ distribution. The Premier League data and R code to reproduce the illustration of Section 5 are available in the public repository https://github.com/crisvarin/peb.